# thanks: These authors contribute equally to this workthanks: These authors contribute equally to this workthanks: These authors contribute equally to this work

### v2

@@ Abstract
The emergence of the arrow of time in quantum many-body systems stems from the inherent tendency of Hamiltonian evolution to scramble quantum information and increase entanglement. While, in principle, one might counteract this temporal directionality by engineering a perfectly inverted Hamiltonian to reverse entanglement growth, such a scenario is fundamentally unstable because even minor imperfections in the backward evolution can be exponentially amplified, a hallmark of quantum many-body chaos. Therefore, successfully reversing quantum many-body dynamics demands a deep understanding of the underlying structure of quantum information scrambling and chaotic dynamics. Here, by using solid-state nuclear magnetic resonance on a macroscopic ensemble of randomly interacting spins, we measure the out-of-time-ordered correlator (OTOC) and validate key predictions of scramblon theory, a universal theoretical framework for information scrambling. Crucially, this theory enables us to isolate and mitigate errors in the OTOC caused by imperfections in the backward evolution. As a result, this protocol uncovers the anticipated exponential behavior of quantum many-body chaos and extracts the quantum Lyapunov exponent in a many-body experimental system for the first time. Our results push the fundamental limits of dynamical reversibility of complex quantum systems, with implications for quantum simulation and metrology.
@@
Quantum information scrambling lies at the heart of thermalization in isolated quantum systems [1, 2, 3, 4, 5]. When a low-entanglement initial state evolves under a Hamiltonian $\hat{H}$, local information spreads across the entire system, becoming irretrievable via local measurements. At the same time, the entanglement entropy grows, rendering the system locally a thermal state. Intriguingly, one can envision a gedanken experiment: if we apply the reversed Hamiltonian $-\hat{H}$ to this thermal state, is it possible to recover the original low-entanglement state and its local information? This question echoes a profound analogy in black hole physics: whether information thrown in and thermalized by a black hole can ever be retrieved. Now, advances in quantum control techniques on various synthetic quantum platforms attempt to turn such a thought experiment into a tangible ambition [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31], raising an even deeper question: can engineered inverted dynamics defy the arrow of time?
@@
In realistic quantum systems, however, such an oversimplified scenario does not materialize. As illustrated by Fig. 1a, even infinitesimal imperfections in the engineered reversed Hamiltonian can lead to an exponentially amplified error as evolution time increases, a hallmark of quantum many-body chaos [32, 33, 34, 35, 36]. Given that no practical implementation can perfectly realize $-\hat{H}$, the inherent chaotic nature of quantum many-body dynamics fundamentally prohibits the retrieval of information from a thermal state. The same reasoning extends analogously to the challenge of recovering information from a black hole. Recent experimental observations, spanning nuclear magnetic resonance (NMR) systems [6, 7, 8, 9, 10, 11, 12, 13, 14], trapped ions [15, 16, 17, 18], superconducting qubits [19, 20, 21, 22, 23], ultracold atoms [24, 25, 26, 27, 28, 29, 30], and nitrogen vacancy centers [31], have also directly observed the inevitable effects of imperfect backward evolution.
@@
Figure 1: Schematic of Experimental Protocol. (a) Chaotic quantum many-body dynamics: Unitary evolution under $\hat{H}$ scrambles quantum information, transforming a low-entanglement initial state into a highly entangled state. Under time-reversed dynamics with perturbed Hamiltonian $-\hat{H}+\delta\hat{H}$, the system exhibits exponential deviation from perfect state recovery. (b) Microscopic structure of adamantane (C10H16) powder: Each granule contains adamantane molecules arranged in a face-centered cubic lattice. Individual molecules consist of spin-$1/2$ 1H nuclei and spinless 12C atoms. (c) Top: Floquet pulse sequence converting the dipolar Hamiltonian Eq. (1) into the engineered form Eq. (2). Bottom: Experimental sequence for MQC and OTOC measurements.
@@
In this work, we measure the OTOC in a solid-state NMR sample with a macroscopic number of randomly interacting spins, a system beyond the capabilities of classical simulation. This can be achieved using a well-established protocol for measuring the multiple quantum coherence (MQC) spectrum (Fig. 1c, bottom panel) [43, 44]. The measurement involves the reverse evolution of this many-body system, which, of course, is imperfect. We utilize scramblon theory to derive a fitting ansatz for the measured OTOC, achieving good agreement with the experiment. Moreover, the scramblon theory imposes several strong constraints on the fitting parameters, and we verify that these constraints are well satisfied by the fitting results. In other words, the experimental data strongly validate the scramblon theory. With this fitting formula, we can extrapolate the measured OTOC to the error-free limit. We demonstrate that the OTOC after this error mitigation exhibits the anticipated long-time behavior, in sharp contrast to other methods explored in the previous literature.
@@
The experiment is performed on a powdered adamantane (C10H16) sample, as illustrated in Fig. 1b and detailed in the supplementary material [45]. Each adamantane molecule contains 16 nuclear spin-$1/2$ hydrogen nuclei, denoted by $\mathbf{S}_{im}$, where the index $i$ labels the molecules and $m=1,\dots,16$ identifies individual spins within a molecule. Due to rapid molecular thermal motion, intramolecular spin interactions are averaged out, leaving only the intermolecular dipolar coupling as the dominant spin-spin interaction. The sample is placed in a uniform $9.4$ T magnetic field, which aligns the spin quantization axes and induces fast Larmor precession of all spins. In the rotating frame and under the secular approximation, the effective dipolar Hamiltonian takes the form [46]:
@@
Then, by applying periodic global radio-frequency pulse sequences to rotating spins, as depicted in Fig. 1c, a more general effective Hamiltonian can be engineered via the Floquet-Magnus expansion as $\hat{H}_{\text{eff}}=\hat{H}_{0}+\hat{H}_{1}$. $\hat{H}_{0}$ represents the leading-order contribution in the expansion, which can be generally written as:
@@
where $\mu,\nu=x,y,z$. The protocol generating Eq. (2) imposes specific constraints on the coefficients $\xi_{\mu\nu}$, namely the traceless condition $\sum_{\mu=x,y,z}\xi_{\mu\mu}=0$ and the symmetry condition $\xi_{\mu\nu}=\xi_{\nu\mu}$. An important observation is that if a set $\{\xi_{\mu\nu}\}$ satisfies these constraints, so does the set $\{-\xi_{\mu\nu}\}$. This implies that Floquet engineering always enables reversal of $\hat{H}_{0}$. However, the higher-order terms, denoted by $\hat{H}_{1}$, generally cannot be simultaneously reversed. Therefore, without loss of generality, we denote the forward evolution as governed by $\hat{H}=\hat{H}_{0}+\hat{H}_{1}$, and the backward evolution as governed by $-\hat{H}_{0}-\hat{H}_{1}^{\prime}=-\hat{H}-\delta\hat{H}$. Here, $\delta\hat{H}$, representing the difference between $\hat{H}_{1}$ and $\hat{H}^{\prime}_{1}$, is the imperfection in the reversed dynamics. In this system, $\hat{H}_{1}$ and $\delta\hat{H}$ are typically on the order of a few percent relative to $\hat{H}_{0}$ [45, 14].
The measurement protocol is schematically illustrated in Fig. 1(c). We begin by defining the total spin operator along the $\gamma$-axis as $\hat{O}_{\gamma}=\sum_{im}\hat{S}^{\gamma}_{im}$, where $\gamma\in\{x,y,z\}$. The system is initialized in a weakly polarized state described by the density matrix $\hat{\rho}_{\gamma}\propto\mathbbm{1}+\epsilon\hat{O}_{\gamma}$, with $\epsilon\ll 1$ characterizing the small polarization. Following forward time evolution under the Hamiltonian $\hat{H}$, the density matrix transforms as $\hat{\rho}_{\gamma}(t)={\rm e}^{-{\rm i}\hat{H}t}\hat{\rho}_{\gamma}{\rm e}^{{\rm i}\hat{H}t}$. Subsequently, we implement a uniform spin rotation by an angle $\phi$ around the $\alpha$-axis ($\alpha\in\{x,y,z\}$), yielding the transformed density matrix: $\hat{\rho}_{\alpha\gamma}(\phi,t)={\rm e}^{-{\rm i}\hat{O}_{\alpha}\phi}{\rm e}^{-{\rm i}\hat{H}t}\hat{\rho}_{\gamma}{\rm e}^{{\rm i}\hat{H}t}{\rm e}^{{\rm i}\hat{O}_{\alpha}\phi}$. This is followed by backward evolution as described above. Finally, we measure the total magnetization along the $\gamma$-direction $\langle\hat{O}_{\gamma}\rangle$, which equals the correlation function $F_{\alpha\gamma}(\phi,t)$ defined as:
@@
Here, $\hat{\mathcal{V}}(t)\equiv\exp\left(-{\rm i}\mathbf{T}\int_{0}^{t}{\rm d}t^{\prime}\delta\hat{H}(t^{\prime})\right)$ represents the time-evolution operator in the interaction picture, with $\mathbf{T}$ denoting time ordering. All operators $\hat{O}(t)$ and $\delta\hat{H}(t)$ evolve according to the Heisenberg picture under $\hat{H}$. For normalization, we divide $F_{\alpha\gamma}(\phi,t)$ by $\mathcal{C}=\langle\hat{O}_{\gamma}^{2}\rangle=N/4$ to ensure $F_{\alpha\gamma}(0,0)=1$. The MQC spectrum can be obtained through the Fourier transformation of $F_{\alpha\gamma}(\phi,t)$ with respect to $\phi$ [43, 44].
We now highlight the relationship between $F_{\alpha\gamma}(\phi,t)$ and OTOCs through two limits: (i) For perfect reversed dynamics ($\delta\hat{H}=0$), $F_{\alpha\gamma}(\phi,t)$ reduces to
@@
and this OTOC, also known as the Loschmidt echo, probes imperfections in the reversed dynamics. Both cases can be computed by scramblon theory [41], strongly suggesting that the general behavior of $F_{\alpha\gamma}(\phi,t)$, although more challenging to derive, can also be captured by scramblon theory. Indeed, as we show in the Supplementary Material [45], an ansatz for $F_{\alpha\gamma}(\phi,t)$ in the small-$\phi$ regime can be derived as:
@@
where $\{a,b,\varkappa,\Delta\}$ are all fitting parameters. The $\phi=0$ limit of Eq. (6) reproduces the Loschmidt echo (Eq. (5)), with $a\neq 0$ quantifying the reversal imperfection. We note that a recent theoretical work also discussed a similar issue in a specific solvable model [47].
Figure 3: Verifying the Predictions of Scramblon Theory. The three columns correspond to the fitting parameters extracted from three different Hamiltonians: $\hat{H}^{(1)}$, $\hat{H}^{(2)}$, and $\hat{H}^{(3)}$. Subfigures a-c in the upper row display the fitting parameters for different values of $\alpha$ and $\gamma$, with green diamonds representing $a\times 50$, yellow squares representing $\Delta$, and blue triangles representing $\varkappa$. The dashed lines represent $\varkappa$ averaged over $\alpha$ and $\gamma$, and $a$ and $\Delta$ averaged over $\alpha$. Subfigures d-f in the lower row show the eigenvalues $\sigma_{1},\sigma_{2},\sigma_{3}$ for the symmetric matrix $M^{\text{s}}$ and the absolute value of eigenvalues $\sigma_{4},\sigma_{5},\sigma_{6}$ for the antisymmetric matrix $M^{\text{a}}$. The error bars on the fitting parameters represent the $95\%$ confidence interval, including the standard deviation from each fitting residual $\sigma_{\text{res}}$ and the uncertainty arising from the choice of fitting range $\sigma_{\text{range}}$ [45].
We investigate three representative chaotic Hamiltonians: (i) $\hat{H}_{0}^{(1)}$ with $\xi_{xx}=-\xi_{yy}=0.4$, (ii) $\hat{H}_{0}^{(2)}$ with $\xi_{xy}=\xi_{yx}=0.4$, and (iii) $\hat{H}_{0}^{(3)}$ with $(\xi_{xx},\xi_{yy},\xi_{zz})=(0.2,-0.225,0.025)$, while all other $\xi_{\mu\nu}=0$. For each Hamiltonian $\hat{H}^{(\lambda)}_{0}$, we perform measurements for all nine possible combinations of $\alpha,\gamma\in\{x,y,z\}$ axes, yielding a total of $3\times 9=27$ distinct cases. For each case, we measure three different $\phi$ angles and fit the three curves jointly with a single scramblon ansatz, Eq. (6). We find that all data are excellently captured by the scramblon ansatz Eq. (6), with three representative fits shown in Fig. 2 and the complete set of all data available in the Supplementary Material [45].
@@
Fig. 3 demonstrates how we verify these three constraints for our three Hamiltonians. Panels (a-c) display the values of $a$, $\Delta$, and $\varkappa$ for all nine measurement combinations per Hamiltonian, with dashed lines indicating (i) $\varkappa$ averaged over $\gamma$ and $\alpha$, and (ii) $\alpha$-averaged $a$ and $\Delta$ values that depend only on $\gamma$. Most cases show only a few percent deviations from averaged values, with a maximum deviation of $\sim 16\%$ occasionally, confirming the first two constraints. To test the third constraint, we decompose $M$ into symmetric ($M^{\text{s}}=(M+M^{\text{T}})/2$) and antisymmetric ($M^{\text{a}}=(M-M^{\text{T}})/2$) components. Eigenvalues shown in panels (d-f) reveal that $M^{\text{s}}$ exhibits only one dominant eigenvalue, and all eigenvalues of $M^{\text{a}}$ are negligible, validating the rank-1 symmetric requirement for $M$.
@@
The excellent agreement with experimental data strongly validates the scramblon ansatz. Within this framework, the parameter $a$ fully captures imperfections in the time-reversed dynamics, as evidenced by $F_{\alpha\gamma}(0,t)\equiv 1$ when $a=0$. This allows an error mitigation by setting $a=0$ in Eq. (6) while keeping other fitted parameters unchanged (blue curve in Fig. 4a). For comparison, we examine an alternative mitigation approach from previous literatures that normalizes $F_{\alpha\gamma}(\phi,t)$ by $F(0,t)$ to define $\tilde{F}_{\alpha\gamma}(\phi,t)=F_{\alpha\gamma}(\phi,t)/F_{\alpha\gamma}(0,t)$ (red triangles in Fig. 4a) [19, 15, 9, 48]. Within the scramblon formalism, this yields:
@@
Different approaches are compared in Fig. 4a.
@@
which for perfect time reversal ($\delta\hat{H}=0$) corresponds to the OTO commutator $\langle|[\hat{O}_{\alpha}(t),\hat{O}_{\gamma}]|^{2}\rangle$ and exhibits exponential growth, reflecting the rapid increase of the operator size $\hat{O}_{\alpha}(t)$ under Heisenberg evolution before reaching the scrambling time. Analysis based on the scramblon ansatz reveals: (i) The raw data (Eq. (6)) gives $I_{\alpha\gamma}(t)=2b\Delta{\rm e}^{\varkappa t}/(1+a{\rm e}^{\varkappa t})^{1+\Delta}$, which decays at late times due to $\Delta>0$ (green dashed line in Fig. 4b); (ii) The normalized approach (Eq. (7)) yields $I_{\alpha\gamma}(t)=2b\Delta{\rm e}^{\varkappa t}/(1+a{\rm e}^{\varkappa t})$, saturating to $2b\Delta/a$ at long times (red dash-dotted line in Fig. 4b); and (iii) Only our $a=0$ mitigation produces the physically expected exponential growth $2b\Delta{\rm e}^{\varkappa t}$ (blue solid line in Fig. 4b), where $\varkappa$ emerges as the quantum Lyapunov exponent. These distinct behaviors are quantitatively compared with experimental data. To the best of our knowledge, $\varkappa$ shown in Fig. 3 serves as the first exact experimental determination of the Lyapunov exponent in a quantum many-body system with macroscopic degrees of freedom.
@@
In conclusion, we develop an ansatz based on scramblon theory to describe reversed quantum many-body dynamics in the presence of imperfections. We validate this ansatz by fitting it to experimental data from an NMR system with macroscopically randomly interacting spins, where the stringent constraints imposed by scramblon theory are convincingly verified by the fitting outputs. This approach allows us to mitigate the effects of imperfections and recover error-resilient reversed dynamics. Notably, time-reversal-based exponential sensitivity has also been recently exploited to enhance quantum metrology in systems such as cold atoms in cavities [24, 25] and nitrogen vacancy centers [31], where imperfections also suppress this sensitivity at long evolution times. Our theory may be extended to such quantum metrology applications, potentially enabling further improvements in sensitivity [49].

### v3

@@ Abstract
The emergence of the arrow of time in quantum many-body systems stems from the inherent tendency of Hamiltonian evolution to scramble quantum information and increase entanglement. While, in principle, one might counteract this temporal directionality by engineering a perfectly inverted Hamiltonian to reverse entanglement growth, such a scenario is fundamentally unstable because even minor imperfections in the backward evolution can be exponentially amplified, a hallmark of quantum many-body chaos. Therefore, successfully reversing quantum many-body dynamics demands a deep understanding of the underlying structure of quantum information scrambling and chaotic dynamics. In this letter, by using solid-state nuclear magnetic resonance on a macroscopic ensemble of randomly interacting spins, we measure the out-of-time-ordered correlator (OTOC) and validate key predictions of scramblon theory, a universal theoretical framework for information scrambling. Crucially, this theory enables us to isolate and mitigate errors in the OTOC caused by imperfections in the backward evolution. As a result, this protocol uncovers the anticipated exponential behavior of quantum many-body chaos and extracts the quantum Lyapunov exponent in a many-body experimental system for the first time. Our results push the fundamental limits of dynamical reversibility of complex quantum systems, with implications for quantum simulation and metrology.
@@
Quantum information scrambling lies at the heart of thermalization in isolated quantum systems [1, 2, 3, 4, 5]. When a low-entanglement initial state evolves under a Hamiltonian $\hat{H}$, local information spreads across the entire system, becoming irretrievable via local measurements. At the same time, the entanglement entropy grows, rendering the system locally a thermal state. Intriguingly, one can envision a gedanken experiment: if we apply the reversed Hamiltonian $\hat{H}$ to this thermal state, is it possible to recover the original low-entanglement state and its local information? This question echoes a profound analogy in black hole physics: whether information thrown in and thermalized by a black hole can ever be retrieved. Now, advances in quantum control techniques on various synthetic quantum platforms attempt to turn such a thought experiment into a tangible ambition [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31], raising an even deeper question: can engineered inverted dynamics defy the arrow of time?
@@
In realistic quantum systems, however, such an oversimplified scenario does not materialize. As illustrated by Fig. 1(a), even infinitesimal imperfections in the engineered reversed Hamiltonian can lead to an exponentially amplified error as evolution time increases, a hallmark of quantum many-body chaos [32, 33, 34, 35, 36]. Given that no practical implementation can perfectly realize $\hat{H}$, the inherent chaotic nature of quantum many-body dynamics fundamentally prohibits the retrieval of information from a thermal state. The same reasoning extends analogously to the challenge of recovering information from a black hole. Recent experimental observations, spanning nuclear magnetic resonance (NMR) systems [6, 7, 8, 9, 10, 11, 12, 13, 14], trapped ions [15, 16, 17, 18], superconducting qubits [19, 20, 21, 22, 23], ultracold atoms [24, 25, 26, 27, 28, 29, 30], and nitrogen vacancy centers [31], have also directly observed the inevitable effects of imperfect backward evolution.
@@
Figure 1: Schematic of Experimental Protocol. (a) Chaotic quantum many-body dynamics: Unitary evolution under $\hat{H}$ scrambles quantum information, transforming a low-entanglement initial state into a highly entangled state. Under time-reversed dynamics with perturbed Hamiltonian $\hat{H}+\delta\hat{H}$, the system exhibits exponential deviation from perfect state recovery. (b) Microscopic structure of adamantane (C10H16) powder: Each granule contains adamantane molecules arranged in a face-centered cubic lattice. Individual molecules consist of spin-$1/2$ 1H nuclei and spinless 12C atoms. (c) Top: Floquet pulse sequence converting the dipolar Hamiltonian Eq. (1) into the engineered form Eq. (2) [37, 8]. Bottom: Experimental sequence for MQC and OTOC measurements.
@@
In this work, we measure the OTOC in a solid-state NMR sample with a macroscopic number of randomly interacting spins, a system beyond the capabilities of classical simulation. This can be achieved using a well-established protocol for measuring the multiple quantum coherence (MQC) spectrum (Fig. 1(c), bottom panel) [44, 45]. The measurement involves the reverse evolution of this many-body system, which, of course, is imperfect. We utilize scramblon theory to derive a fitting ansatz for the measured OTOC, achieving good agreement with the experiment. Moreover, the scramblon theory imposes several strong constraints on the fitting parameters, and we verify that these constraints are well satisfied by the fitting results. In other words, the experimental data strongly validate the scramblon theory. With this fitting formula, we can extrapolate the measured OTOC to the error-free limit. We demonstrate that the OTOC after this error mitigation exhibits the anticipated long-time behavior, in sharp contrast to other methods explored in the previous literature[19, 15, 9].
@@
Physical system and experimental protocol—The experiment is performed on a powdered adamantane (C10H16) sample, as illustrated in Fig. 1(b) and detailed in the supplementary material [46]. Each adamantane molecule contains 16 nuclear spin-$1/2$ hydrogen nuclei, denoted by $\mathbf{S}_{im}$, where the index $i$ labels the molecules and $m=1,\dots,16$ identifies individual spins within a molecule. Due to rapid molecular thermal motion, intramolecular spin interactions are averaged out, leaving only the intermolecular dipolar coupling as the dominant spin-spin interaction. The sample is placed in a uniform $9.4$ T magnetic field, which aligns the spin quantization axes and induces fast Larmor precession of all spins. In the rotating frame and under the secular approximation, the effective dipolar Hamiltonian takes the form [66]:
@@
Then, by applying periodic global radio-frequency pulse sequences to rotating spins, as depicted in Fig. 1(c), a more general effective Hamiltonian can be engineered via the Floquet-Magnus expansion as $\hat{H}_{\text{eff}}=\hat{H}_{0}+\hat{H}_{1}$. $\hat{H}_{0}$ represents the leading-order contribution in the expansion, which can be generally written as:
@@
where $\mu,\nu=x,y,z$. The protocol generating Eq. (2) imposes specific constraints on the coefficients $\xi_{\mu\nu}$, namely the traceless condition $\sum_{\mu=x,y,z}\xi_{\mu\mu}=0$ and the symmetry condition $\xi_{\mu\nu}=\xi_{\nu\mu}$. An important observation is that if a set $\{\xi_{\mu\nu}\}$ satisfies these constraints, so does the set $\{-\xi_{\mu\nu}\}$. This implies that Floquet engineering always enables reversal of $\hat{H}_{0}$. However, the higher-order terms, denoted by $\hat{H}_{1}$, generally cannot be simultaneously reversed. Therefore, without loss of generality, we denote the forward evolution as governed by $\hat{H}=\hat{H}_{0}+\hat{H}_{1}$, and the backward evolution as governed by $\hat{H}_{0}-\hat{H}_{1}^{\prime}=-\hat{H}-\delta\hat{H}$. Here, $\delta\hat{H}$, representing the difference between $\hat{H}_{1}$ and $\hat{H}^{\prime}_{1}$, is the imperfection in the reversed dynamics. In this system, $\hat{H}_{1}$ and $\delta\hat{H}$ are typically on the order of a few percent relative to $\hat{H}_{0}$ [46, 14].
The measurement protocol is schematically illustrated in Fig. 1(c). We begin by defining the total spin operator along the $\gamma$-axis as $\hat{O}_{\gamma}=\sum_{im}\hat{S}^{\gamma}_{im}$, where $\gamma\in\{x,y,z\}$. The system is initialized in a weakly polarized state described by the density matrix $\hat{\rho}_{\gamma}\propto\mathbbm{1}+\epsilon\hat{O}_{\gamma}$, with $\epsilon\ll 1$ characterizing the small polarization. Following forward time evolution under the Hamiltonian $\hat{H}$, the density matrix transforms as $\hat{\rho}_{\gamma}(t)={\rm e}^{-{\rm i}\hat{H}t}\hat{\rho}_{\gamma}{\rm e}^{{\rm i}\hat{H}t}$. Subsequently, we implement a uniform spin rotation by an angle $\phi$ around the $\alpha$-axis ($\alpha\in\{x,y,z\}$), yielding the transformed density matrix: $\hat{\rho}_{\alpha\gamma}(\phi,t)={\rm e}^{-{\rm i}\hat{O}_{\alpha}\phi}{\rm e}^{-{\rm i}\hat{H}t}\hat{\rho}_{\gamma}{\rm e}^{{\rm i}\hat{H}t}{\rm e}^{{\rm i}\hat{O}_{\alpha}\phi}$. This is followed by backward evolution as described above. Finally, we measure the total magnetization along the $\gamma$-direction $\langle\hat{O}_{\gamma}\rangle$, which equals the correlation function $\epsilon F_{\alpha\gamma}(\phi,t)$ defined as:
@@
Here, $\hat{\mathcal{V}}(t)\equiv\mathbf{T}\exp\left(-{\rm i}\int_{0}^{t}{\rm d}t^{\prime}\delta\hat{H}(t^{\prime})\right)$ represents the time-evolution operator in the interaction picture, with $\mathbf{T}$ denoting time ordering. All operators $\hat{O}(t)$ and $\delta\hat{H}(t)$ evolve according to the Heisenberg picture under $\hat{H}$. For normalization, we divide $F_{\alpha\gamma}(\phi,t)$ by $\mathcal{C}=\langle\hat{O}_{\gamma}^{2}\rangle=N/4$ to ensure $F_{\alpha\gamma}(0,0)=1$. The MQC spectrum can be obtained through the Fourier transformation of $F_{\alpha\gamma}(\phi,t)$ with respect to $\phi$ [44, 45].
Validating the scramblon theory—We now highlight the relationship between $F_{\alpha\gamma}(\phi,t)$ and OTOCs through two limits: (i) For perfect reversed dynamics ($\delta\hat{H}=0$), $F_{\alpha\gamma}(\phi,t)$ reduces to
@@
and this OTOC, also known as the Loschmidt echo, probes imperfections in the reversed dynamics. Both cases can be computed by scramblon theory [42, 67], strongly suggesting that the general behavior of $F_{\alpha\gamma}(\phi,t)$, although more challenging to derive, can also be captured by scramblon theory. Indeed, as we show in the Supplementary Material [46], an ansatz for $F_{\alpha\gamma}(\phi,t)$ in the small-$\phi$ regime can be derived as:
@@
where $\{a,b,\varkappa,\Delta\}$ are all fitting parameters that depend on $\alpha$ and $\gamma$. The $\phi=0$ limit of Eq. (6) reproduces the Loschmidt echo (Eq. (5)), with $a\neq 0$ quantifying the reversal imperfection. We note that a recent theoretical work also discussed a similar issue in a specific solvable model [68].
Figure 3: Verifying the Predictions of Scramblon Theory. The three columns correspond to the fitting parameters extracted from three different Hamiltonians: $\hat{H}^{(1)}$, $\hat{H}^{(2)}$, and $\hat{H}^{(3)}$. Subfigures (a)-(c) in the upper row display the fitting parameters for different values of $\alpha$ and $\gamma$, with green diamonds representing $a\times 50$, yellow squares representing $\Delta$, and blue triangles representing $\varkappa$. The dashed lines represent $\varkappa$ averaged over $\alpha$ and $\gamma$, and $a$ and $\Delta$ averaged over $\alpha$. Subfigures (d)-(f) in the lower row show the eigenvalues $\sigma_{1},\sigma_{2},\sigma_{3}$ for the symmetric matrix $M^{\text{s}}$ and the absolute value of eigenvalues $\sigma_{4},\sigma_{5},\sigma_{6}$ for the antisymmetric matrix $M^{\text{a}}$. The error bars on the fitting parameters represent the $95\%$ confidence interval, including the standard deviation from each fitting residual $\sigma_{\text{res}}$ and the uncertainty arising from the choice of fitting range $\sigma_{\text{range}}$ [46].
We investigate three representative chaotic Hamiltonians: (i) $\hat{H}_{0}^{(1)}$ with $\xi_{xx}=-\xi_{yy}=0.4$, (ii) $\hat{H}_{0}^{(2)}$ with $\xi_{xy}=\xi_{yx}=0.4$, and (iii) $\hat{H}_{0}^{(3)}$ with $(\xi_{xx},\xi_{yy},\xi_{zz})={(0.2,-0.225,0.025)}$, while all other $\xi_{\mu\nu}=0$. For each Hamiltonian $\hat{H}^{(\lambda)}_{0}$, we perform measurements for all nine possible combinations of $\alpha,\gamma\in\{x,y,z\}$ axes, yielding a total of $3\times 9=27$ distinct cases. For each case, we measure three different $\phi$ angles and fit the three curves jointly with a single scramblon ansatz, Eq. (6). We find that all data are excellently captured by the scramblon ansatz Eq. (6), with three representative fits shown in Fig. 2. We quantify the goodness of fit using the adjusted $R^{2}$, as detailed in the Supplementary Material [46]. Discrepancies between experiment and theory may arise from the invalidity of the scramblon theory at very early times $e^{\varkappa t}\lesssim 1$.
@@
Fig. 3 demonstrates how we verify these three constraints for our three Hamiltonians. Panels (a-c) display the values of $a$, $\Delta$, and $\varkappa$ for all nine measurement combinations per Hamiltonian, with dashed lines indicating (i) $\varkappa$ averaged over $\gamma$ and $\alpha$, and (ii) $\alpha$-averaged $a$ and $\Delta$ values that depend only on $\gamma$. Most cases show only a few percent deviations from averaged values, with a maximum deviation of $\sim 16\%$ occasionally, confirming the first two constraints. To test the third constraint, we decompose $M$ into symmetric ($M^{\text{s}}=(M+M^{\text{T}})/2$) and antisymmetric ($M^{\text{a}}=(M-M^{\text{T}})/2$) components. Eigenvalues shown in panels (d-f) reveal that $M^{\text{s}}$ exhibits only one dominant eigenvalue, and all eigenvalues of $M^{\text{a}}$ are negligible, validating the rank-1 symmetric requirement for $M$. It is notable that a strong correspondence between the results for $H^{(1)}_{0}$ and $H^{(2)}_{0}$ is observed. This is anticipated, as $H^{(1)}_{0}$ can be transformed into $H^{(2)}_{0}$ via a global rotation ${\rm e}^{-{\rm i}\hat{O}_{z}\pi/4}$.
@@
Recovering exponential behavior of OTOC and extracting the Lyapunov exponent—The excellent agreement with experimental data strongly validates the scramblon ansatz. Within this framework, the parameter $a$ fully captures imperfections in the time-reversed dynamics, as evidenced by $F_{\alpha\gamma}(0,t)\equiv 1$ when $a=0$. This allows an error mitigation by setting $a=0$ in Eq. (6) while keeping other fitted parameters unchanged (blue curve in Fig. 4(a)). For comparison, we examine an alternative mitigation approach from previous literatures that normalizes $F_{\alpha\gamma}(\phi,t)$ by $F(0,t)$ to define $\tilde{F}_{\alpha\gamma}(\phi,t)=F_{\alpha\gamma}(\phi,t)/F_{\alpha\gamma}(0,t)$ (red triangles in Fig. 4(a)) [19, 15, 9, 69]. Within the scramblon formalism, this yields:
@@
Different approaches are compared in Fig. 4(a).
@@
which for perfect time reversal ($\delta\hat{H}=0$) corresponds to the OTO commutator $\langle|[\hat{O}_{\alpha}(t),\hat{O}_{\gamma}]|^{2}\rangle$ and exhibits exponential growth, reflecting the rapid increase of the operator size $\hat{O}_{\alpha}(t)$ under Heisenberg evolution before reaching the scrambling time. Analysis based on the scramblon ansatz reveals: (i) The raw data (Eq. (6)) gives $I_{\alpha\gamma}(t)=4b\Delta{\rm e}^{\varkappa t}/(1+a{\rm e}^{\varkappa t})^{1+2\Delta}$, which decays at late times due to $\Delta>0$ (green dashed line in Fig. 4(b)); (ii) The normalized approach (Eq. (7)) yields $I_{\alpha\gamma}(t)=4b\Delta{\rm e}^{\varkappa t}/(1+a{\rm e}^{\varkappa t})$, saturating to $4b\Delta/a$ at long times (red dash-dotted line in Fig. 4(b)); and (iii) Only our $a=0$ mitigation produces the physically expected exponential growth $4b\Delta{\rm e}^{\varkappa t}$ (blue solid line in Fig. 4(b)), where $\varkappa$ emerges as the quantum Lyapunov exponent. It is also remarkable that only our $a=0$ mitigated result $I_{\alpha\gamma}(t)$ exhibits the symmetric and rank-1 property predicted by the scramblon theory for systems with time-reversal symmetry, because there is a single dominant scramblon mode and its coupling to both $\hat{O}_{\alpha}$ and $\hat{O}_{\gamma}$ operators is symmetric (see Supplementary Materiral [46] for more experimental data on all combinations of $\alpha$ and $\gamma$). These distinct behaviors are quantitatively compared with experimental data. To the best of our knowledge, $\varkappa$ shown in Fig. 3 serves as the first exact experimental determination of the Lyapunov exponent in a quantum many-body system with macroscopic degrees of freedom.
@@
Conclusion—In conclusion, we develop an ansatz based on scramblon theory to describe reversed quantum many-body dynamics in the presence of imperfections. We validate this ansatz by fitting it to experimental data from an NMR system with macroscopically randomly interacting spins, where the stringent constraints imposed by scramblon theory are convincingly verified by the fitting outputs. These results suggest that the applicability of the scramblon theory extends beyond idealized all-to-all interacting models, remaining valid in realistic systems with microscopic long-range interactions. The application of the scramblon theory allows us to mitigate the effects of imperfections and recover error-resilient reversed dynamics. Notably, time-reversal-based exponential sensitivity has also been recently exploited to enhance quantum metrology in systems such as cold atoms in cavities [24, 25] and nitrogen vacancy centers [31], where imperfections also suppress this sensitivity at long evolution times. Our theory may be extended to such quantum metrology applications, potentially enabling further improvements in sensitivity.